OpenStudy (toxicsugar22):

how do u solve (2-x+3x^2)(3+2x-x^2+2x^4)

OpenStudy (amistre64):

whats your definition of solve?

OpenStudy (amistre64):

to me; solve would be to distribute each term in the left thru the paranethesis and then comcine like terms

OpenStudy (toxicsugar22):

sorry Find the first three terms of the product in ascending powers (2-x+3x^2)(3+2x-x^2+2x^4)

OpenStudy (amistre64):

ascending eh...from little power to biggest...

OpenStudy (toxicsugar22):

yes

OpenStudy (amistre64):

6 +4x -3x+9x^2-2x^2-2x^2...if I see it right :) 6 +x +5x^2

OpenStudy (amistre64):

and I think i see others....hold that

OpenStudy (amistre64):

3 +2x -x^2 * 2 -x +3x^2 -------------- 9x^2+0+0 -3x-2x^2+0 6+4x-2x^2 ------------------- 6 +x +5x^2.....yeah, that should be it

OpenStudy (toxicsugar22):

thanks you

OpenStudy (toxicsugar22):

can u help me with one more

OpenStudy (amistre64):

i can try

OpenStudy (toxicsugar22):

and this time let me do it and you do it also and then we can compare ou answers to see if i got it

OpenStudy (amistre64):

k

OpenStudy (toxicsugar22):

(1-2x^2-x^4)(4+x^2-2x^4)

OpenStudy (amistre64):

i got mine :)

OpenStudy (toxicsugar22):

i got mine

OpenStudy (toxicsugar22):

so i got 4+9x^2+5x^6

OpenStudy (amistre64):

4 - 6x^2 -8x^4 ??

OpenStudy (toxicsugar22):

is it right

OpenStudy (amistre64):

the lowest powers are x^0; x^2 and x^4

OpenStudy (amistre64):

4 - 7x^2 -8x^4 ...had a typo :) 4x^0 = 4 x^2 -8x^2 = 7x^2 -2x^4 -2x^2 -4x^4 = -8x^4

OpenStudy (toxicsugar22):

0k - got it

OpenStudy (toxicsugar22):

thanks